OpenStudy (anonymous):

Find the inverse of this function: f(x) = 3 - 2tan(0.5x) I'm not really sure how to do this... I've tried y = 3 - 2tan(0.5x) and rearranging to make x the subject

hartnn (hartnn):

can u first isolate the tan(0.5x) term from y=...?

OpenStudy (anonymous):

-2tan(0.5x)=y-3 tan(0.5x)=(y-3)/(-2)?

hartnn (hartnn):

correct, when tan y = z, then y= arctan z so, tan 0.5x = (y-3)/(-2) 0.5 x = arctan [(y-3)/(-2)] got this ?

OpenStudy (anonymous):

What exactly is arctan?

hartnn (hartnn):

its the inverse tan function. for details, look wiki

OpenStudy (anonymous):

Oh its the inverse, we never learnt that term in new zealand lol. So x = 2 arctan[(y-3)/(-2)]?

hartnn (hartnn):

yes, correct... you can write that as 2 arctan[(3-x)/(2)] also which is your inverse function

OpenStudy (anonymous):

Thanks so much!!! You seem to be the most active person on here lol.

hartnn (hartnn):

there are many others active here too :) and welcome ^_^